OpenStudy (anonymous):

This isn't a math problem but my math final is tomorrow and I need some help. If there are 510 points possible in the assignment course, and 200 points on the final. I have a 385 points in the assignment course giving me a 75%... If I get a 45% on the final, what would my over-all grade be? Thanks!

OpenStudy (anonymous):

.45*200 = 90 pts. 385+90 = 475 475/510 = 93.1

OpenStudy (anonymous):

or maybe my question should be, what grade do i need to get on the final to have a 70% or higher?

OpenStudy (anonymous):

youve already passed the class. 385/510 = 75%

OpenStudy (anonymous):

and i believe the 200 final points will be tacked onto the 510 making it 720... sorry about that

OpenStudy (anonymous):

oh oh, my bad.

OpenStudy (anonymous):

i mean 710** wow i am bad at math! lol

OpenStudy (anonymous):

ok, so we want to solve: \[\frac{385+x}{710} = .70\] to figure out what you need in order to get a 70 in the class

OpenStudy (anonymous):

wow, i don't see how you do it! amazing lol

OpenStudy (anonymous):

this turns into: \[385+x = 497 \Rightarrow x = 497 - 385 = 112\]

OpenStudy (anonymous):

that means id have to make a 112?

OpenStudy (anonymous):

yes, you need to score at least a 112. higher and you pass, lower and you dont. 112/200 = .56 which is 56% So its like saying you need to get a 56 on the test.

OpenStudy (anonymous):

ok, i can probably handle that hopefully, 20 question test, multiple choice! Thank you!